# Empirical Bayes Estimation in Heterogeneous Coefficient Panel Models

### v1

@@ 6.1 WFR Algorithm for the HC Model
In practice, the following procedure can be used to initialize the support points and weights. For $j=1,\ldots,m$:
@@
    Sample indices $i(1),\ldots,i(B)$ uniformly at random from $\{1,\ldots,n\}$.
@@
    as the subsample MLE based on $\{(Y_{i(b)},X_{i(b)})\}_{b=1}^{B}$, and assign a uniform weight $\omega_{0}^{j}=1/m$.
@@ 7 Empirical Analysis of Income Dynamics
There is a large body of literature seeking to understand heterogeneity in earnings. See Altonji, Hynsjö, and Vidangos (2023) and Browning and Ejrnaes (2013) for recent reviews of this literature. We highlight several closely related methodological contributions. Chamberlain and Hirano (1999) consider a parametric Bayesian approach that allow for heterogeneity in $\sigma_{i}^{2}$. Geweke and Keane (2000) also adopt a Bayesian framework with a finite mixture model for the composite errors $\sigma_{i}e_{it}$, while Hirano (2002) extend this framework using an infinite mixture model. Gu and Koenker (2017) propose an EB approach that was closest in spirit to ours. These four papers did not allow for random $(b_{i},\rho_{i})$. Giacomini, Lee, and Sarpietro (2025) develop an individual-weight shrinkage estimator that optimizes unit-level (rather than aggregate) accuracy by exploiting each individual’s own past history, with feasible weights motivated by minimax regret. Moon, Schorfheide, and Zhang (2023) consider random components in all $(a_{i},b_{i},\rho_{i},\sigma_{i}^{2})$ within a parametric Bayesian framework. Specifically, they adopt a spike-and-slab prior that accommodate either dense or sparse structures but assume independence across random components. In contrast, we adopt a nonparametric EB approach for the prior distribution and allow for dependence among the random components.
Gu and Koenker (2017, GK hereafter) analyze log real earnings data studied in Meghir and Pistaferri (2004). The extract consists of 938 individuals who have continuous earnings records from age 25 onward in the Panel Study of Income Dynamics (PSID) for the period 1968-1993. The panel is unbalanced panels with varying numbers of time periods $T_{i}$. [^12] The only minor difference between the current model and the original HIVDX model in (1) is that $T_{i}$ may vary across $i$. The likelihood function can be easily modified to accommodate unbalanced panels. Their $Y_{it}$ denotes residuals obtained from year-specific regressions of log real earnings on a vector of covariates. Though the HIVDX model motivates GK’s analysis, they impose $\rho_{i}=\rho_{*}$ for some unknown homogeneous parameter $\rho_{*}$ and $b_{i}\equiv 0$ to circumvent computational challenges. Furthermore, their likelihood is conditioned on the initial observation $Y_{i1}$, while implicitly assuming that the random component is independent of $Y_{i1}$. Thus, the marginal likelihood depends on $Y_{i1}$ only through $Y_{i2}-\rho_{*}Y_{i1}$.
@@
where the hats indicate either EB or individual-level MLE estimates. The EB estimators are obtained by reestimating the NPMLE using the sample that excludes the last-period observations, i.e., $\{(Y_{it},X_{2,it}):i=1,\ldots,N;t=1,\ldots,T_{i}-1\}$. As shown in Figure 1, both predictions and prediction errors from EB exhibit substantial shrinkage compared to those based on MLE, resulting in variance reductions of approximately $25\%$ for predictions and $20\%$ for prediction errors.
@@
We now discuss the estimation results concerning the joint distribution of the random components. As shown in Table 1, a notable feature is the negative correlation between $a_{i}$ and $b_{i}$. To further examine this relationship, the left panel of Figure 3 plots the EB estimate pairs $(\hat{a}_{i}^{\mathrm{EB}},\hat{b}_{i}^{\mathrm{EB}})$ for $N=938$ individuals, revealing a clear negative association between the individual intercepts $\hat{a}_{i}^{\mathrm{EB}}$ and slopes $\hat{b}_{i}^{\mathrm{EB}}$. [^14] For completeness, Figure 4 in Appendix C presents all pairwise scatter plots of the EB estimates for $\theta_{i}$. To explore the experience profile in the cross-section, we use the estimated moments in Table 1 to compute the cross-sectional prior variance of $a_{i}+b_{i}x$, which captures the deviation of an individual’s income trajectory at $X_{2,it}=x$ from the average. The mapping $x\mapsto\mathbb{V}_{\hat{G}}(a_{i}+b_{i}x)$, shown in the right panel of Figure 3, indicates that the variance declines up to about $x=1$ (ten years of experience) and increases thereafter. This non-monotonic pattern arises from the negative covariance between $a_{i}$ and $b_{i}$. Interestingly, a similar U-shaped relationship between the log variance of residual earnings and experience is documented in Mincer (1974, Chapter 6, p.104), who attributes it to “a weak correlation between post-school investment ratios and earning capacity.”
@@ 8 Monte Carlo Experiments
    The Restricted HIVD model follows the HIVD model above but imposes $\rho_{i}\equiv 0.5$ and assumes $\sigma_{i}^{2}$ has a two-point distribution satisfying $G_{*}(\sigma_{i}^{2}=0.1)=G_{*}(\sigma_{i}^{2}=0.3)=0.5$. The initial distribution of $u_{i0}$ is $\mathcal{N}(0,\sigma_{i}^{2}/(1-0.5^{2}))$. The distribution of $a_{i}$ and the sample size is the same as in HIVD. This scenario is that a researcher does not know that $\rho_{i}$ is a constant and continue to estimate the HIVD model.
@@
    The HIDVX model generates $\theta_{i}=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ from the prior distribution estimated from the PSID sample. The deterministic covariate (interpreted as experience divided by 10) is assumed to evolve according to $X_{2,it}=X_{2,i,t-1}+0.1$, with $(N,T)=(1000,13)$, approximately as in the PSID sample. The initial value $X_{2,i1}$ is drawn from a discrete distribution with probabilities $\mathbb{P}(X_{2,i1}=0.3)=0.25$, $\mathbb{P}(X_{2,i1}=0.4)=0.05$, $\mathbb{P}(X_{2,i1}=0.5)=\mathbb{P}(X_{2,i1}=0.6)=0.1$, and $\mathbb{P}(X_{2,i1}=0.7)=0.5$.
@@
$\mathrm{Bias}=\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i},\qquad\mathrm{SD}=\left(\frac{1}{N}\sum_{i=1}^{N}\left(\hat{\varepsilon}_{i}-\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i}\right)^{2}\right)^{1/2},\qquad\mathrm{RMSE}=\left(\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i}^{2}\right)^{1/2}.$
@@
Table 2 reports results for the HIVD model and its restricted variant. In both designs, the oracle estimator achieves the lowest estimation error, but the EB estimator improves upon the individual MLE with uniformly smaller RMSE across all parameters. The gains are particularly pronounced for $\rho_{i}$. In line with this theoretical prediction, the gains from EB are even larger in the restricted HIVD design, where the prior for $\rho_{i}$ is more informative. Most of the RMSEs of the EB estimator are within 10% of those of the oracle.
@@ 9 Conclusions
We have proposed a nonparametric empirical Bayes framework for analyzing short-panel data with rich forms of unobserved heterogeneity. The analysis is a significant generalization of classical G-modeling to accommodate heterogeneous slopes and non-spherical error structures. Identification, consistency, and optimality results are established under general conditions, with primitive sufficient conditions derived for specific cases of interest. However, we have not developed any accompanying inference methods. Recent contributions such as Armstrong, Kolesár, and Plagborg-Møller (2022) and Ignatiadis and Wager (2022) can be useful for future research.
Appendix A Proofs
A.1 Proofs of Main Results
@@
This equality follows from integrating both sides of (A.1) with respect to $t\in\mathbb{R}^{T-{d_{\beta}}}$, ensuring that both $\tilde{H}_{*}$ and $\tilde{H}_{o}$ are probability measures. We are prepared to apply Lemma 1. Since $\delta\mapsto(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1}$ is injective by Assumption 2(ii), (30) is satisfied by $\Lambda(\mathcal{K}_{\delta})$ which consists only of $(\lambda_{1}(\delta),\lambda_{2}(\delta)):=(0,(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1})$. By Lemma 1, it follows that $\tilde{H}_{*}=\tilde{H}_{o}$, and hence $H_{*}=H_{o}$.
@@
Our next goal is to establish that the real and imaginary parts of $\mu$ are both zero. By construction of $\mu$, (A.1) implies that
@@
where $\mu_{\pm}^{R}$ are mutually singular non-negative Borel measures, representing the positive and negative real parts of $\mu$, respectively. The same holds for the positive and negative imaginary parts $\mu_{\pm}^{I}$. By linearity of $\mathcal{T}$ with respect to $\mu$, it follows from (A.1) that
@@
Again, it suffices to verify Assumption 2. See Appendix B.1 for details. ∎
@@
Case 3(b): Assume that there are consecutive equal terms in $r_{t}$. Without loss of generality, assume that $r_{1}=r_{2}=\kappa$. Substituting $r_{1}$ and $r_{2}$ into (A.1) with $t=1$, the ratio of (A.1) to (25) identifies
@@
$\leq$ $\ \frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{2}|X_{2,iT+1}|^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{\rho}_{i}^{{\mathrm{EB}}}-\rho_{i}^{*})^{2}|Y_{iT}|^{2}+\frac{1}{N}\sum_{i=1}^{N}(\widehat{a\rho}_{i}^{{\mathrm{EB}}}-(a\rho)_{i}^{*})^{2}$
$\quad+\frac{1}{N}\sum_{i=1}^{N}(\widehat{b\rho}_{i}^{{\mathrm{EB}}}-(b\rho)_{i}^{*})^{2}|X_{2,i,T}|^{2}$
$\leq$ $\ \frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{\rho}_{i}^{{\mathrm{EB}}}-\rho_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|Y_{iT}|^{4}}$
@@
$\mathbb{E}\left[\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}}\right]\leq\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}\right]}\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}\right]}.$
@@
A.2 Technical Lemmas
A.2.1 Identification Lemma (Theorem 1)
@@
A.2.2 Consistency Lemmas (Theorem 2)
@@
A.2.3 Regret Consistency Lemmas (Theorem 3, Proposition 5)
@@
Appendix B Auxiliary Identification Results
B.1 Verification of Assumption 2 for ARMA(1,1) Errors
@@
B.2 Identification Failure of the HIVDX model with $T=4$
@@
Appendix C Additional Empirical Results

### v2

@@ 6.1 WFR Algorithm for the HC Model
In practice, we generate a dispersed initialization by computing MLEs on random subsamples of the data. The procedure is as follows. Fix the subsample size $B$. If the time dimension $T$ is sufficient to ensure that individual MLEs are well-defined, one may choose $B=1$ to maximize the diversity of the grid points. For $j=1,\ldots,m$:
@@
    Sample a set of indices $\mathcal{I}_{j}=\{i(1),\ldots,i(B)\}$ uniformly at random from $\{1,\ldots,N\}$.
@@
    and assign a uniform weight $\omega_{0}^{j}=1/m$.
@@ 7 Empirical Analysis of Income Dynamics
There is a large body of literature seeking to understand heterogeneity in earnings. See Altonji, Hynsjö, and Vidangos (2023) and Browning and Ejrnaes (2013) for recent reviews of this literature. We highlight several closely related methodological contributions. Chamberlain and Hirano (1999) consider a parametric Bayesian approach that allow for heterogeneity in $\sigma_{i}^{2}$. Geweke and Keane (2000) also adopt a Bayesian framework with a finite mixture model for the composite errors $\sigma_{i}e_{it}$, while Hirano (2002) extends this framework using an infinite mixture model. Gu and Koenker (2017) propose an EB approach that was closest in spirit to ours. These four papers did not allow for random $(b_{i},\rho_{i})$. Giacomini, Lee, and Sarpietro (2025) develop an individual-weight shrinkage estimator that optimizes unit-level (rather than aggregate) accuracy by exploiting each individual’s own past history, with feasible weights motivated by minimax regret. Moon, Schorfheide, and Zhang (2023) consider random components in all $(a_{i},b_{i},\rho_{i},\sigma_{i}^{2})$ within a parametric Bayesian framework. Specifically, they adopt a spike-and-slab prior that accommodate either dense or sparse structures but assume independence across random components. In contrast, we adopt a nonparametric EB approach for the prior distribution and allow for dependence among the random components.
In this section, we re-examine the application in Gu and Koenker (2017, GK hereafter), which analyzes log real earnings data studied in Meghir and Pistaferri (2004). The extract consists of 938 individuals who have continuous earnings records from age 25 onward in the Panel Study of Income Dynamics (PSID) for the period 1968–1993. The panel is unbalanced with varying numbers of time periods $T_{i}$. [^11] The only minor difference between the current model and the original HIVDX model in (1) is that $T_{i}$ may vary across $i$. The likelihood function can be easily modified to accommodate unbalanced panels. As in Gu and Koenker (2017), $Y_{it}$ denotes residuals obtained from year-specific regressions of log real earnings on a vector of covariates. Though the HIVDX model motivates GK’s analysis, they assume $\rho_{i}$ is homogeneous and $b_{i}=0$ in the application. Furthermore, their likelihood, expressed in terms of sufficient statistics, is conditioned on the initial observation $Y_{i1}$, which implicitly assumes that the random component is independent of $Y_{i1}$. Thus, their marginal likelihood depends on $Y_{i1}$ only through $Y_{i2}-\rho_{*}Y_{i1}$.
@@
where the hats indicate either EB or individual-level MLE estimates. As shown in Figure 1, the EB method exhibits substantial shrinkage, reducing the variance of the predicted values by approximately 25% compared to the MLE. This translates into improved forecasting performance, leading to a variance reduction of approximately 20% in the prediction errors.
@@
As shown in Table 1, a notable feature is the negative correlation between $a_{i}$ and $b_{i}$. To further examine this relationship, the left panel of Figure 3 plots the EB estimate pairs $(\hat{a}_{i}^{\mathrm{EB}},\hat{b}_{i}^{\mathrm{EB}})$ for $N=938$ individuals, revealing a clear negative association between the individual intercepts $\hat{a}_{i}^{\mathrm{EB}}$ and slopes $\hat{b}_{i}^{\mathrm{EB}}$. [^13] Figure 4 in Appendix O-2 presents all pairwise scatter plots of the EB estimates for $\theta_{i}$. To explore the experience profile in the cross-section, we use the estimated moments in Table 1 to compute the cross-sectional prior variance of $a_{i}+b_{i}x$, which captures the deviation of an individual’s income trajectory at $X_{2,it}=x$ from the average. The mapping $x\mapsto\mathbb{V}_{\hat{G}}(a_{i}+b_{i}x)$, shown in the right panel of Figure 3, indicates that the variance declines up to about $x=1$ (ten years of experience) and increases thereafter. This non-monotonic pattern arises from the negative covariance between $a_{i}$ and $b_{i}$. Interestingly, a similar U-shaped relationship between the log variance of residual earnings and experience is documented in Mincer (1974, Chapter 6, p.104), who attributes it to “a weak correlation between post-school investment ratios and earning capacity.”
@@
For completeness, we also computed estimates using the EM algorithm discussed in Section 6.2, initializing it with the same starting values used for WFR (as described in Section 6.1). We find that the EM estimates are very similar to the WFR solution. While the EM algorithm attains a slightly higher likelihood value by optimizing weights over the dense initial grid, the WFR algorithm yields an approximate solution with fewer mass points. Furthermore, although the EM algorithm generally requires fewer iterations to converge, the computational cost per iteration is higher. These results are encouraging, as they suggest WFR is well-suited for higher-dimensional problems where fixed-grid algorithms become computationally challenging.
@@ 8 Monte Carlo Experiments
    The Restricted HIVD model follows the HIVD model above but imposes $\rho_{i}\equiv 0.5$ and assumes $\sigma_{i}^{2}$ has a two-point distribution satisfying $G_{*}(\sigma_{i}^{2}=0.1)=G_{*}(\sigma_{i}^{2}=0.3)=0.5$. The initial distribution of $u_{i0}$ is $\mathcal{N}(0,\sigma_{i}^{2}/(1-0.5^{2}))$. The distribution of $a_{i}$ and the sample size is the same as in HIVD. This DGP mimics a scenario in which a researcher does not know that $\rho_{i}$ is a constant and continue to estimate the HIVD model.
@@
    The HIDVX model generates $\theta_{i}=(a_{i},b_{i},\sigma_{i}^{2},\rho_{i})$ from the prior distribution estimated from the PSID sample. The covariate (interpreted as experience divided by 10) is assumed to evolve according to $X_{2,it}=X_{2,i,t-1}+0.1$, with $(N,T)=(1000,13)$, approximately as in the PSID sample. The initial value $X_{2,i1}$ is drawn from a discrete distribution with probabilities $\mathbb{P}(X_{2,i1}=0.3)=0.25$, $\mathbb{P}(X_{2,i1}=0.4)=0.05$, $\mathbb{P}(X_{2,i1}=0.5)=\mathbb{P}(X_{2,i1}=0.6)=0.1$, and $\mathbb{P}(X_{2,i1}=0.7)=0.5$.
@@
$\mathrm{Bias}=\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i},\quad\mathrm{SD}=\left(\frac{1}{N}\sum_{i=1}^{N}\left(\hat{\varepsilon}_{i}-\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i}\right)^{2}\right)^{1/2},\quad\mathrm{RMSE}=\left(\frac{1}{N}\sum_{i=1}^{N}\hat{\varepsilon}_{i}^{2}\right)^{1/2}.$
@@
Table 2 reports results for the HIVD model and its restricted variant. In both designs, the oracle estimator achieves the lowest estimation error, but the EB estimator improves upon the individual MLE with uniformly smaller RMSE across all parameters. The gains are particularly pronounced for $\rho_{i}$. The gains from EB are even larger in the restricted HIVD design, where the prior for $\rho_{i}$ is more informative. Most of the RMSEs of the EB estimator are within 10% of those of the oracle.
@@ 9 Conclusions
We have proposed a nonparametric empirical Bayes framework for analyzing short-panel data with rich forms of unobserved heterogeneity. The analysis generalizes classical G-modeling to accommodate heterogeneous slopes and non-spherical error structures. Identification, consistency, and optimality results are established under general conditions, with primitive sufficient conditions derived for specific cases of interest. However, we have not developed any accompanying inference methods. Recent contributions such as Armstrong, Kolesár, and Plagborg-Møller (2022) and Ignatiadis and Wager (2022) can be useful for future research.
Supplement to “Empirical Bayes Estimation in Heterogeneous Coefficient Panel Models”
Appendix S-1 Proofs
S-1.1 Proofs of Main Results
@@
This equality follows from integrating both sides of (S-1.1) with respect to $t\in\mathbb{R}^{T-{d_{\beta}}}$, ensuring that both $\tilde{H}_{*}$ and $\tilde{H}_{o}$ are probability measures. We are prepared to apply Lemma 1. Since $\delta\mapsto(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1}$ is injective by Assumption 2(ii), (29) is satisfied by $\Lambda(\mathcal{K}_{\delta})$ which consists only of $(\lambda_{1}(\delta),\lambda_{2}(\delta)):=(0,(\boldsymbol{M}^{\prime}P(\delta)P(\delta)^{\prime}\boldsymbol{M})^{-1})$. By Lemma 1, it follows that $\tilde{H}_{*}=\tilde{H}_{o}$, and hence $H_{*}=H_{o}$.
@@
Our next goal is to establish that the real and imaginary parts of $\mu$ are both zero. By construction of $\mu$, (S-1.1) implies that
@@
where $\mu_{\pm}^{R}$ are mutually singular non-negative Borel measures, representing the positive and negative real parts of $\mu$, respectively. The same holds for the positive and negative imaginary parts $\mu_{\pm}^{I}$. By linearity of $\mathcal{T}$ with respect to $\mu$, it follows from (S-1.1) that
@@
Again, it suffices to verify Assumption 2. See Appendix O-1. ∎
@@
Case 3(b): Assume that there are consecutive equal terms in $r_{t}$. Without loss of generality, assume that $r_{1}=r_{2}=\kappa$. Substituting $r_{1}$ and $r_{2}$ into (S-1.1) with $t=1$, the ratio of (S-1.1) to (24) identifies
@@
$\leq$ $\ \frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{2}|X_{2,iT+1}|^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{\rho}_{i}^{{\mathrm{EB}}}-\rho_{i}^{*})^{2}|Y_{iT}|^{2}$
$\quad+\frac{1}{N}\sum_{i=1}^{N}(\widehat{a\rho}_{i}^{{\mathrm{EB}}}-(a\rho)_{i}^{*})^{2}+\frac{1}{N}\sum_{i=1}^{N}(\widehat{b\rho}_{i}^{{\mathrm{EB}}}-(b\rho)_{i}^{*})^{2}|X_{2,i,T}|^{2}$
$\leq$ $\ \frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}^{{\mathrm{EB}}}-a_{i}^{*})^{2}+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}}$
$\quad+\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{\rho}_{i}^{{\mathrm{EB}}}-\rho_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|Y_{iT}|^{4}}$
@@
$\mathbb{E}\left[\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}}\right]\\ \leq\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}(\hat{b}_{i}^{{\mathrm{EB}}}-b_{i}^{*})^{4}\right]}\sqrt{\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}|X_{2,iT+1}|^{4}\right]}.$
@@
S-1.2 Technical Lemmas
S-1.2.1 Identification Lemma (Theorem 1)
@@
S-1.2.2 Consistency Lemmas (Theorem 2)
@@
S-1.2.3 Regret Consistency Lemmas (Theorem 3, Proposition 5)
@@
Appendix O-1 Auxiliary Identification Results
O-1.1 Verification of Assumption 2 for ARMA(1,1) Errors
@@
O-1.2 Identification Failure of the HIVDX model with $T=4$
@@
Appendix O-2 Additional Empirical Results